# Object Agnostic 3D Lifting in Space and Time

### v1

@@ Abstract
We present a spatio-temporal perspective on category-agnostic 3D lifting of 2D keypoints over a temporal sequence. Our approach differs from existing state-of-the-art methods that are either: (i) object agnostic, but can only operate on individual frames, or (ii) can model space-time dependencies, but are only designed to work with a single object category. Our approach is grounded in two core principles. First, when there is a lack of data about an object, general information from similar objects can be leveraged for better performance. Second, while temporal information is important, the most critical information is in immediate temporal proximity. These two principles allow us to outperform current state-of-the-art methods on per-frame and per-sequence metrics for a variety of objects. Lastly, we release a new synthetic dataset containing 3D skeletons and motion sequences of a diverse set animals. Dataset and code will be made publicly available.
Figure 0: Left: Bottom row shows the predicted 3D skeletons of a puma category in motion. The blue sticks represent our model’s predictions, closely tracking the red ground-truth positions, showing our model’s ability to generate smooth and precise motion over time. The dashed line highlights the trajectory of a specific joint $\hat{\mathbf{Y}}_{t,j}$, emphasizing the temporal consistency and accuracy of our approach. Right: Quantitative SA-MPJPE comparison across 13 animal categories, where our method consistently outperforms competing models.
@@ 1 Introduction
Reconstructing 3D deforming objects from 2D landmarks is a long-standing challenge in computer vision. While traditional approaches use multiple camera views to mitigate ambiguities, recent learning-based methods are capable of lifting 3D object structure using sparse 2D keypoints from a single camera. Coupled with an abundance of publicly available 3D human pose data, human-specific lifting models like MotionBERT [34] have become increasingly capable. However, their reliance on human-specific information and vast amounts of training data make it problematic if they are to be used for other objects. In particular, animals pose a significant challenge due to the limited amount of publicly available 3D animal data.
This has motivated recent developments around object-agnostic lifting. Most notably, 3D-LFM [7] achieved state-of-the-art lifting performance by leveraging a combined dataset of various object categories. It leverages the permutation equivariant property of transformers, along with additional inductive biases, to handle object category imbalance and within-category object rig/skeleton variation. However, a shortfall of 3D-LFM is its inability to utilize temporal information, leading to less accurate reconstructions in dynamic sequences.
Our work tackles these challenges by introducing the first object-agnostic 3D lifting framework that is both data-efficient and temporally aware. Our approach uses the power of transformers with a strategic inductive bias that focuses on temporally proximate frames, enabling it to capture motion dynamics effectively. This approach improves the accuracy of 3D reconstructions across various object categories, even in scenarios with limited data and unseen objects. Additionally, we address the lack of publicly available datasets for diverse animal skeletons and motion sequences by creating a new synthetic dataset. This dataset includes 4D labels for 13 animal categories, encompassing 698 animation sequences with temporal consistency, designed to foster further research in class-agnostic lifting.
@@
  • We propose a class-agnostic lifting model with a strategic inductive bias directly embedded in the architecture. We demonstrate the state-of-the-art 3D lifting of our approach across scenarios involving noise, occlusions, and unseen objects.
  • We contribute a new synthetic dataset containing 4D skeletons for a variety of animals with animated behaviour sequences, where temporal consistency is prioritised through a non-linear refinement procedure.
We empirically validate the effectiveness of our approach on our synthetic dataset. We achieve state-of-the-art results with existing metrics and provide an additional metric to provide a more complete analysis. We plan to release all dataset and code.
@@ 2 Related Works
Obtaining the 3D pose of an object from a single monocular camera generally follows one of two paths. The first approach directly predicts the 3D pose from RGB images [2, 25, 22, 26], but this method often struggles with generalization to distribution shifts, such as varying lighting and background conditions. Alternatively, a two-stage method divides the task between two specialized models [11, 3, 20, 4]: a pose detector first extracts the 2D pose, which is then lifted into 3D by a separate model. Our work aligns with this two-stage approach, focusing on improving the robustness and generalization of the 3D lifting step.
@@ 2.2 Object-Specific Lifting
Traditional non-rigid structure-from-motion (NRSfM) algorithms have been effective for modeling simple objects, such as human bodies and hands [5, 9]. These methods rely on the availability of 2D keypoints and specific 3D supervision for the object in question. However, recent deep learning approaches have demonstrated superior performance in handling the complexity of various object rigs [24, 13, 28, 17]. Despite these advancements, they still require the 2D keypoints to have consistent semantic correspondence across all instances of the object, meaning that a specific landmark, such as an elbow, must have the same semantic meaning across different poses.
This limitation persists even in state-of-the-art deep learning models like MotionBERT [34] and others [3, 4], which are tailored specifically for human body 3D lifting. The specialized nature of these models and their dependency on large datasets make them unsuitable for objects with limited available data, such as animals. Animal-specific 3D lifting models suffer from similar issues, often being restricted to a single animal category and demonstrating poor generalization due to data scarcity [15, 21, 10].
@@ 2.3 Object-Agnostic Lifting
The paradigm of object-agnostic lifting has been recently pioneered by 3D-LFM [7], which can handle a wide range of object categories by leveraging large-scale information of large-scale data to enhance performance for underrepresented or unseen objects. However, unlike object-specific models like Motionbert [34], 3D-LFM does not incorporate temporal information, which is needed for accurate 3D reconstruction over sequences. Our framework builds upon this insight, integrating the strengths of both object-agnostic lifting and temporal modeling to achieve improved results.
@@ 2.4 Animal Datasets
To effectively benchmark category-agnostic lifting over videos of animals, a diverse set of animal categories with accurately labeled 3D skeletons is essential. Recent animal datasets have made strides in this area but are often limited to single atemporal images [29, 31], with few publicly available datasets offering 3D poses of animals in video sequences [18, 6]. Moreover, methods to collect data for a broad set of animals are typically impractical, expensive, or yield noisy results.
Inspired by the tracking community, we turn to synthetic data [8, 14, 33]. Existing synthetic datasets are often restricted to a single animal, such as pigs [1] or ants [27], and generally contain only simplistic motion sequences. The DeformableThings4D [19] dataset offers more complex and diverse motion sequences animated by artists, but it was created for dense mesh recovery and does not include 3D skeletons. We build upon these models and animations to create a new dataset specifically designed for the task of class-agnostic 3D lifting.
@@ 3 Method
Figure 1: Overview of our data pipeline and 3D lifting model. The left side of the figure demonstrates (a) the process of calculating joints from animal mesh vertices, and (b) the projection of the skeleton joints into 2D keypoints. The right side of the figure illustrates our lifting model at a high-level. The sequence of 2D input and temporal index is projected and passed through our motion encoder and space encoder layers. The spatio-temporal latent features are decoded into canonical 3D structures. The canonical structures are then aligned to the ground truth (GT) via procrustes-alignment for calculating the loss.
@@ 3.1 Dataset
We find a noticeable gap in available public datasets containing diverse 3D animal skeletons with realistic motion sequences. We aim to fill this gap by creating a new synthetic dataset building on the mesh vertices and animation sequences provided by the DeformableThings4D [19] dataset. We provide 3D skeleton labels for 13 animal categories, totalling 698 animation sequences with temporal correspondence across 51,360 frames. We provide statistics and examples of our dataset in the supp. material. We detail our data collection pipeline in following section, see Fig. 1 for a visualization of our data collection pipeline.
@@ 3.1.1 Animal joints
Given the vertices of a skin-tight mesh of an animal and an associated sequence of movement, our aim is to find the 3D locations of anatomically-accurate skeleton joints of the animal. We define the locations of $K$ vertices and $N$ joints in 3D space throughout a sequence of $T$ frames as $\mathbf{V}\in\mathbb{R}^{T\times K\times 3}$ and $\mathbf{J}\in\mathbb{R}^{T\times N\times 3}$, respectively. Note that the number of mesh vertices, joints, and frames may vary across animals and sequences. We take inspiration from motion capture, where markers are attached to an object and used to estimate joint positions via triangulation. Let us define $j\in\{1,2,...,N\}$. We strategically select a subset of vertices $\mathbf{S}_{t,j}=\bigl{\{}\mathbf{S}_{t,j,1},...,\mathbf{S}_{t,j,m}\bigl{\}}\subset\mathbf{V}_{t}$ to be virtual markers, such that the mean of the virtual markers will provide the location of a joint $j$ in frame $t$:
@@
The selection of vertices for a subset $\mathbf{S}_{t,j}$ is guided by a visual inspection of $\mathbf{J}_{t,j}$ and the motion similarity of vertices within $\mathbf{S}_{t,j}$. The bones of the animal are defined as an adjacency matrix $\mathbf{A}^{N\times N}$ containing the connections between joints. We decide the number of joints, their approximate locations, and their connections by reviewing the anatomical structure of the target animal.
@@ 3.1.2 Non-linear optimisation
The noise that is inherently present in our human annotation process occasionally results in the length of animal bones to change over time, measured as the Euclidean distance between two connected joints. We adopt an additional inverse-kinematics optimisation to refine the position of joints such that the bone lengths are consistent. We formulate the problem as solving for the pose angles $\theta\in\mathbb{R}^{T\times N\times 3}$ of forward kinematics for each joint $j$ in frame $t$:
@@
We use gradient descent to optimise the objective function
@@
Our inverse-kinematics optimisation provides a new set of joints $\mathbf{\widehat{J}}_{t,j}$ ensuring that an animal has consistent bone lengths throughout any complex sequence of movement.
@@ 3.1.3 Perspective projection
We define an initial camera pose to satisfy two conditions. The mean location of the animal throughout the sequence is at the center of the camera view, and all joints are within view for the entirety of the sequence. We randomly rotate the camera around the y-axis and then project the points to the 2D camera plane. For our purpose we do this only once for each animation sequence, however this process can be used to obtain many different views of the animal throughout the sequence.
@@ 3.2 Lifting model
Given an input sequence of 2D skeletons $\mathbf{X}\in\mathbb{R}^{T\times J\times 2}$, where $T$ is the number of frames in the video and $J$ the number of joints, our goal is to reconstruct the 3D skeletons $\mathbf{\widehat{Y}}\in\mathbb{R}^{T\times J\times 3}$ of the object for each frame.
@@ 3.2.1 Keypoint features
The attention mechanism of transformers is inherently permutation equivariant, such that inputs can be randomly permuted and the corresponding outputs will remain the same. We leverage this property to handle objects with different joint configurations. We utilise the masking mechanism of [7] to handle the technical challenge of training with different numbers of joints. Inputs are zero-padded up to the maximum number of joints in a mini-batch and a mask $\mathbf{M}\in\{0,1\}^{J}$ is used to ignore padded joints. Each element in $\mathbf{M}$ is defined as:
@@
We project the 2D skeletons to $D$-dimensional features $\mathbf{F}\in\mathbb{R}^{T\times J\times D}$ using Random Fourier Features (RFF) [32]. We additionally encode each 2D joint $(x,y)$ with its temporal location $t$ in the video. We compute the feature of an input $\mathbf{p}=[x,y,t]^{T}$ as:
@@
where $\mathbf{W}\in\mathbb{R}^{\frac{D}{2}\times 3}$ is sampled from a normal distribution $\mathcal{N}(0,I)$ and $\mathbf{b}\in\mathbb{R}^{\frac{D}{2}}$ is sampled from a uniform distribution $\mathcal{U}(0,\frac{1}{2\pi})$. We choose analytical RFF for its success in low-data and out-of-distribution (OOD) scenarios [7, 32]. We find that encoding the temporal position is beneficial for our motion encoder.
@@ 3.2.2 Motion encoder
Our motion encoder leverages multi-head self-attention (MHSA) to embed temporal context into the keypoint features. Although we choose MHSA for its ability to use information from all frames, its lack of any inductive bias makes it unsuitable for tasks with limited data. We argue that it is not necessary to consider all frames in a sequence because most of the useful information can be found in nearby frames. We explicitly impose this inductive bias into the MHSA blocks by applying a binary mask to the intermediate attention maps. We define the mask $\mathbf{Z}\in\mathbb{R}^{T\times T}$ for a joint at time $t$ as:
@@
where $\alpha$ is a hyper-parameter controlling the number of frames before and after $t$ that can contribute information during attention..
@@
where $\mathbf{Q}^{(h)},\mathbf{K}^{(h)},\mathbf{V}^{(h)}\in\mathbb{R}^{N,T,\frac{D}{H}}$ for a total of $H$ heads. We apply our temporal mask $\mathbf{Z}$ before the non-linear softmax operation for each head:
@@
$\text{MHSA}_{\alpha}(\mathbf{F}_{M})=[head_{1};...;head_{h}]\mathbf{W_{P}}.$ (10)

### v2

@@ Abstract
We present a spatio-temporal perspective on category-agnostic 3D lifting of 2D keypoints over a temporal sequence. Our approach differs from existing state-of-the-art methods that are either: (i) object-agnostic, but can only operate on individual frames, or (ii) can model space-time dependencies, but are only designed to work with a single object category. Our approach is grounded in two core principles. First, general information about similar objects can be leveraged to achieve better performance when there is little object-specific training data. Second, a temporally-proximate context window is advantageous for achieving consistency throughout a sequence. These two principles allow us to outperform current state-of-the-art methods on per-frame and per-sequence metrics for a variety of animal categories. Lastly, we release a new synthetic dataset containing 3D skeletons and motion sequences for a variety of animal categories.
Figure 1: Left: Bottom row shows the 3D skeletons of a puma animal in motion. The blue lines represent our model’s predictions, closely tracking the red ground-truth lines, demonstrating our model’s ability to generate smooth and precise motion over time. The dashed line highlights the trajectory of a specific joint $\hat{\mathbf{Y}}_{t,j}$, emphasizing the temporal consistency and accuracy of our approach. Right: Quantitative FA-MPJPE comparison across 13 animal categories, where our method consistently outperforms competing models.
@@ 1 Introduction
Reconstructing 3D deforming objects from 2D landmarks obtained by a single camera is a long-standing challenge in computer vision. Traditional non-rigid structure-from-motion (NRSfM) approaches relied on clever but straightforward factorisation methods that are sensitive to noise and occlusions [29, 3]. In some cases, ambiguities can be resolved with multiple camera views, but come at the cost of expensive equipment and limited practicability to natural scenes. Recent learning-based methods are capable of robustly recovering 3D object structure from a single camera in the presence of noise and occlusions. Coupled with an abundance of publicly available 3D human pose data, human-specific lifting models like MotionBERT [36] have become increasingly capable. However, their reliance on human-specific information and vast amounts of training data make it problematic if they are to be used for other objects. In particular, animals pose a significant challenge due to the limited amount of publicly available 3D animal data.
This has motivated recent developments around object-agnostic lifting, where a single model is capable of lifting various object categories without category-specific fine-tuning. Most notably, 3D-LFM [8] achieved state-of-the-art performance on a combined dataset of various object categories. The permutation equivariant property of transformers and additional skeletal information were leveraged to robustly handle category imbalances and within-category object rig/skeleton variations. However, its inability to utilize temporal information results in poor performance when applied to a sequence of 2D poses obtained from a video. We observe frequent jitter and poor recovery of occluded points, particularly in dynamic sequences.
Our work tackles these challenges by introducing the first object-agnostic 3D lifting framework that is both data-efficient and temporally aware. Our approach leverages the power of transformers with a strategic inductive bias that focuses attention on temporally proximate frames, enabling it to effectively capture motion dynamics. Our approach improves the accuracy of 3D reconstructions across various object categories, particularly in challenging scenarios that contain occlusions, fast movement, limited data, and previously unseen categories. Furthermore, we address the lack of publicly available datasets for lifting diverse animal skeletons and motion sequences by creating a new synthetic dataset. Our dataset, AnimalSyn3D, includes 4D labels for 13 animal categories, encompassing 678 animation sequences with temporal consistency, designed to enable further research in class-agnostic lifting.
@@
  • We propose a class-agnostic lifting model with a strategic inductive bias directly embedded in the architecture. We validate the state-of-the-art 3D lifting performance of our approach across challenging scenarios involving noise, occlusions, and unseen objects.
  • We contribute a new synthetic dataset containing 4D skeletons for a variety of animals with animated behavior sequences, where temporal consistency is prioritized through a non-linear refinement procedure.
We empirically validate the effectiveness of our approach on our synthetic dataset. We achieve state-of-the-art results with existing metrics and provide an additional metric for a more complete analysis.
@@ 2 Related Works
Obtaining the 3D pose of an object from a single monocular camera generally follows one of two paths. The first directly predicts the 3D pose from RGB images [2, 26, 23, 27], often struggling to generalize to distribution shifts such as lighting and background information. Alternatively, two-stage methods divide the task between two specialized models [12, 4, 21, 5]: a pose detector first extracts a 2D pose which is then lifted into 3D by a separate model. Our work aligns with a two-stage approach, particularly focusing on improving the robustness and generalization of the 3D lifting stage.
@@ 2.2 Object-Specific Lifting
Traditional NRSfM algorithms have been effective in modeling simple and targeted objects, such as human bodies and hands [6, 10]. These methods largely rely on the availability of 2D keypoints and specific 3D supervision for the object in question. However, recent deep learning approaches have demonstrated superior performance in handling the complexity of various object rigs [25, 14, 30, 18]. Despite these advancements, they still require the 2D keypoints to have consistent semantic correspondence across all instances of the object, where a specific landmark, such as an elbow, must have the same semantic meaning across different poses.
This limitation persists even in state-of-the-art deep lifting models like MotionBERT [36] and others [4, 5], which are tailored specifically for the human body. The specialized nature of these models and their dependence on large datasets make them unsuitable for objects with limited available data, such as animals. Existing animal-specific lifting models suffer from similar issues, often being restricted to a single animal category and demonstrating poor generalization due to data scarcity [16, 22, 11].
@@ 2.3 Object-Agnostic Lifting
The paradigm of object-agnostic lifting has recently been pioneered by 3D-LFM [8], which can handle a wide range of object categories by leveraging large-scale data to enhance performance for underrepresented or unseen objects. However, unlike object-specific models such as Motionbert [36], 3D-LFM does not incorporate temporal information, which is important for accurate 3D reconstruction of sequences. Our framework builds upon this insight, integrating the strengths of both object-agnostic lifting and temporal modeling.
@@ 2.4 Animal Datasets
To effectively benchmark category-agnostic lifting over videos of animals, a diverse set of animal categories with accurately labeled 3D skeletons is required. Recent animal datasets have made strides in this area but are often limited to single atemporal images [31, 33], with few publicly available datasets offering 3D poses of animals in video sequences [19, 7]. Moreover, methods to collect data for a broad set of animals are typically impractical, expensive, or yield noisy results.
Inspired by the tracking community, we turn to synthetic data [9, 15, 35]. Existing synthetic datasets are restricted to a single animal, such as pigs [1] or ants [28], and generally contain only simplistic motion sequences. The DeformableThings4D [20] dataset offers more complex and diverse motion sequences animated by artists, but it was created for dense mesh recovery and does not include 3D skeletons. We build upon these models and animations to create a new dataset specifically designed for the task of class-agnostic 3D lifting.
@@ 3 Method
Figure 2: Overview of our data pipeline and 3D lifting model. The left side of the figure demonstrates (a) the process of calculating skeleton joints from animal mesh vertices, and (b) the projection of the those joints into 2D keypoints. The right side of the figure illustrates our lifting model at a high-level. The sequence of 2D input and temporal index is projected and passed through our motion encoder and space encoder layers. The spatio-temporal latent features are decoded into canonical 3D structures. The canonical structures are then aligned to the ground truth (GT) via procrustes-alignment for calculating the loss.
@@ 3.1 Dataset
We find a noticeable gap in available public datasets containing diverse 3D animal skeletons with realistic motion sequences. We aim to fill this gap by creating a new synthetic dataset, AnimalSyn3D, that builds on the mesh vertices and animation sequences provided by the DeformableThings4D [20] dataset. We provide 3D skeleton labels for 13 animal categories, totalling 678 animation sequences with temporal correspondence across 48,384 frames. We provide statistics and examples of our dataset in the Tab. 7. We detail our data collection pipeline in following sections.
@@ 3.1.1 Animal joints
Given the vertices of a skin-tight mesh of an animal and an associated sequence of movement, our aim is to find the 3D locations of the anatomically-accurate skeleton joints of the animal. We define the locations of $K$ vertices and $N$ joints in 3D space throughout a sequence of $T$ frames as $\mathbf{V}\in\mathbb{R}^{T\times K\times 3}$ and $\mathbf{J}\in\mathbb{R}^{T\times N\times 3}$, respectively. Note that the number of mesh vertices, joints, and frames may vary across animals and sequences. We take inspiration from motion capture, where markers are attached to an object and used to estimate joint positions via triangulation. We strategically select a subset of $M$ vertices $\mathbf{S}_{t,j}=\bigl{\{}\mathbf{S}_{t,j,1},...,\mathbf{S}_{t,j,M}\bigl{\}}\subset\mathbf{V}_{t}$ to be virtual markers, such that the mean of the markers will provide the location of a joint $j$ in frame $t$:
@@
The selection of vertices for a subset $\mathbf{S}_{t,j}$ is guided by a visual inspection of $\mathbf{J}_{t,j}$ and the trajectory similarity of the chosen vertices. The bones of the animal are defined as an adjacency matrix $\mathbf{A}^{N\times N}$ containing the connections between joints. We decide the number of joints, their approximate locations, and their connections by reviewing the anatomical structure of the target animal.
@@ 3.1.2 Non-linear optimisation
The noise that is inherently present in our human annotation process occasionally results in the length of animal bones to change over time. We adopt an additional inverse-kinematics optimization procedure to refine the position of joints so that the bone lengths are consistent across time. We formulate the problem as solving for the pose angles $\theta\in\mathbb{R}^{T\times N\times 3}$ of forward kinematics for each joint $j$ in frame $t$:
@@
We use gradient descent with the Adam [17] optimizer to optimise the objective function
@@
The inverse-kinematics optimization provides a new set of joints $\mathbf{\widehat{J}}_{t,j}$ that ensures consistent bone lengths throughout any complex sequence of movement.
@@ 3.1.3 Perspective projection
We define an initial camera pose to satisfy two conditions. The mean location of the animal throughout the sequence is at the center of the camera view, and all joints are within view for the entirety of the sequence. We randomly rotate the camera around the y-axis and project the points to the 2D camera plane. For our purpose we do this only once for each animation sequence, however this process can be used to obtain many different views of the animal throughout the sequence.
@@ 3.2 Lifting model
Given an input sequence of 2D skeletons $\mathbf{X}\in\mathbb{R}^{T\times J\times 2}$, where $T$ is the number of frames in the video and $J$ is the number of joints, our goal is to reconstruct the 3D skeletons $\mathbf{\widehat{Y}}\in\mathbb{R}^{T\times J\times 3}$ of the object.
@@ 3.2.1 Keypoint features
The attention mechanism of transformers is inherently permutation equivariant, such that inputs can be randomly permuted and the corresponding outputs will remain the same. We leverage this property to handle objects with different joint configurations. We utilise the masking mechanism of [8] to overcome the technical challenge of training with different numbers of joints. Inputs are zero-padded up to the maximum number of joints in a mini-batch and a mask $\mathbf{M}\in\{0,1\}^{J}$ is used to ignore padded joints. Each element in $\mathbf{M}$ is defined as:
@@
We encode the 2D skeletons into $D$-dimensional features $\mathbf{F}\in\mathbb{R}^{T\times J\times D}$ using Random Fourier Features (RFF) [34]. We additionally encode each 2D joint $(x,y)$ with its temporal location $t$ in the sequence. We thus compute the feature of an input $\mathbf{p}=[x,y,t]^{T}$ as:
@@
where $\mathbf{W}\in\mathbb{R}^{\frac{D}{2}\times 3}$ is sampled from a normal distribution $\mathcal{N}(0,I)$ and $\mathbf{b}\in\mathbb{R}^{\frac{D}{2}}$ is sampled from a uniform distribution $\mathcal{U}(0,\frac{1}{2\pi})$. We choose analytical RFF for its success in low-data and out-of-distribution (OOD) scenarios [8, 34]. We find that encoding the temporal position is beneficial for the motion encoder in capturing temporal dependencies.
@@ 3.2.2 Motion encoder
The motion encoder leverages multi-head self-attention (MHSA) to embed temporal context into the keypoint features. Although we choose MHSA for its ability to use information from all frames, its lack of an inductive bias makes it unsuitable for tasks with limited data. We argue that it is not necessary to consider all frames in a sequence because most of the useful information can be found in nearby frames. We explicitly impose this inductive bias into the MHSA blocks by applying a binary mask to the intermediate attention maps. We define the mask $\mathbf{Z}\in\mathbb{R}^{T\times T}$ for a joint at time $t$ as:
@@
where $\alpha$ is a hyper-parameter controlling the number of frames before and after $t$ that can contribute information during attention.
@@
where $\mathbf{Q}^{(h)},\mathbf{K}^{(h)},\mathbf{V}^{(h)}\in\mathbb{R}^{N,T,\frac{D}{H}}$ for a total of $H$ heads. We apply our temporal mask $\mathbf{Z}$ before the non-linear softmax operation in each head:
@@
$\mathbf{F}_{M}=\text{MHSA}_{\alpha}(\mathbf{F}_{M})=[head_{1};...;head_{h}]\mathbf{W_{P}}.$ (10)